How many ways can you arrange 12 books in groups of 3
Answer by stanbon(75887)

How many ways can you arrange 12 books in groups of 3
-----
Pick the groups of 3 in 12C3 = (12*11*10)/(1*2*3) = 220 ways
Arrange the 4 groups 4! = 24 ways
----
Arrange each of the 4 groups in 3! = 6 ways
----
Total # of arrangements: 6*24*220 = 31680
===============
Cheers,
Stan H.
===============